KyledaGreat:

Consider the following relation. 4y+ √x = 4x + (x−1)^2 Evaluate the function found in the previous step at x = 1. f(1) =

KyledaGreat:

@vocaloid

surjithayer:

y=f(1) put x=1 and find the value of y.

KyledaGreat:

y=f(1) = 4y+ √x = 4x + (x−1)^2 you mean ?

surjithayer:

value of y at x=1 is f(1)

surjithayer:

first substitute x=1 and find 4y,then y

KyledaGreat:

f(1) = x=1

surjithayer:

no,put x=1 \[4y+\sqrt{1}=4\times1+(1-1)^2\] 4y=? y=?

KyledaGreat:

y = x - 1/4

surjithayer:

4y+1=4+0 4y=4-1=3 4y=3 y=3/4

surjithayer:

f(1)=value of y at x=1 is 3/4